OpenStudy (anonymous):

If D varies directly with t and D is 105 when t is 3, find D when t is 8. 35 280 840 whahwhawhawhawhhwhattttt

OpenStudy (anonymous):

280

OpenStudy (anonymous):

The formula for direct variation is D = kt.

OpenStudy (anonymous):

First find k.

OpenStudy (anonymous):

105 = k(3) k = 105 / 3 = 35 D = 35t When t = 8, D = 35(8) = 280